### v1

# The Curious Case of Arbitrariness in Machine Learning

Abstract.
Algorithmic modelling relies on limited information in data to extrapolate outcomes for unseen scenarios, often embedding an element of arbitrariness in its decisions. A perspective on this arbitrariness that has recently gained interest is multiplicity—the study of arbitrariness across a set of “good models”, i.e., those likely to be deployed in practice. In this work, we systemize the literature on multiplicity by: (a) formalizing the terminology around model design choices and their contribution to arbitrariness, (b) expanding the definition of multiplicity to incorporate underrepresented forms beyond just predictions and explanations, (c) clarifying the distinction between multiplicity and other traditional lenses of arbitrariness, i.e., uncertainty and variance, and (d) distilling the benefits and potential risks of multiplicity into overarching trends, situating it within the broader landscape of responsible AI. We conclude by identifying open research questions and highlighting emerging trends in this young but rapidly growing area of research.
isbn: 978-1-4503-XXXX-X/18/06copyright: none
1. Introduction
Machine learning attempts to approximate the complexities of the world, inevitably simplifying or generalizing aspects of reality and failing to fully capture its nuances (Singh, 1993; Box, 1976; Barocas et al., 2023; Hooker, 2021; Buolamwini and Gebru, 2018; Birhane, 2022). It is thus inherently susceptible to arbitrariness, as it attempts to extrapolate outcomes based on limited information. Whether due to imperfect data (Buolamwini and Gebru, 2018; Wang et al., 2024; Geirhos et al., 2020), flawed modelling assumptions (Hooker, 2021; Breiman, 2001; Jacobs and Wallach, 2021), or the unpredictability of certain tasks (Wang et al., 2024; Rescher, 1997; Dressel and Farid, 2018), this arbitrariness is an unavoidable byproduct of any data-driven learning, including machine learning. Hence, recognizing and understanding this arbitrariness is crucial for developing responsible learning models.
The study of arbitrariness is not new; it has long been a subject of interest in uncertainty literature, with roots going back centuries in statistics and decision theory (Savage, 1972; de Moivre, 1718; Bayes, 1763; Box and Tiao, 1973; Breiman, 2001; Gal, 2016; Kendall and Gal, 2017). Recently, however, a new paradigm called multiplicity has emerged (Breiman, 2001; Marx et al., 2020; Black et al., 2022). First articulated by Breiman (2001), multiplicity has gained popularity due to its unique focus only on the arbitrariness present within a set of “good models”, i.e., models that pass certain selection criteria and thus are likely to be deployed, commonly known as the Rashomon set. Moreover, multiplicity takes an intriguing perspective on the question of arbitrariness in model decisions, by instead examining arbitrariness in model selection. Through choices made during development and model selection, multiplicity offers an operational lens to the issue of arbitrariness and lays the groundwork for practical solutions in real-world applications.
Several existing works in the literature have provided broad overviews of the field of multiplicity. Black et al. (2022) holds a special place in modern multiplicity research, offering a comprehensive discussion of “opportunities”, “concerns”, and potential “solutions” of multiplicity. Similarly, Rudin et al. (2024) presents an excellent discussion on the benefits of multiplicity, with a focus on its role in identifying simpler and more interpretable models. At this point, it would be remiss not to acknowledge the PhD dissertations of Semenova (2024); Zhong (2024); Cooper (2024); Hsu (2023); Black (2022); Watson-Daniels (2024); Hasan (2022), contributing valuable perspectives on the role of multiplicity in machine learning. Despite these notable contributions, the field still lacks a systematic review of the literature–clearly needed given its rapid growth in recent years (Figure 1). To address this gap, we present the first systematic literature review of multiplicity in machine learning, consolidating existing discussions and identifying overarching trends.
To ensure comprehensive coverage of the literature, we search across various online repositories (DBLP & ACM Digital Library) using multiple search terms (‘rashomon’, ‘model multiplicity’, ‘set of good models’), followed by rigorous manual filtering. We were eventually left with 80 papers that deeply engaged with multiplicity as a central theme in their contributions. Each paper was then manually tagged with all applicable tags, and some relevant statistics are presented in Figure 1. The growing interest in the field is evident, with literature on a wide range of problems related to multiplicity. Precise details about the review process are delegated to the Appendix (§A).
Building on the insights from our review, we make the following contributions to the multiplicity literature. First, we revisit the Rashomon effect, emphasizing the role of developer choices and formalizing the terminology to trace the emergence of various subdomains within multiplicity (§2). Expanding on this discussion, we extend the existing definitions of Rashomon sets and multiplicity to include underrepresented works, making both metrics of model selection and multiplicity more intentional (§3). Next, we distinguish multiplicity from related concepts such as uncertainty and variance; and provide both mathematically grounded differences as well as practical guidance on when to adopt each perspective (§4). Finally, we explore two overarching trends in the multiplicity literature: its role in exploring diverse interpretations during model selection (§5), and its broader implications within responsible AI (§6). We conclude by identifying open research questions and future directions to encourage continued exploration in this domain.
¡2020 2020 2021 2022 2023 2024 001010101020202020303030304040404050505050Year g Number of Papers
BM&E HM REML RSE RAI App Sur PM EM OM 001010101020202020303030304040404050505050Tags
Exploring Alternative InterpretationsResponsible AIMultiplicityRashomon EffectOthers
Figure 1. Statistics from the systematic review, showing the number of papers over the years and their categorization. Each paper can have multiple tags, marking all categories of contributions made by the paper. Details of the tags are provided in the Appendix (§A).
BM&E: Better Models and Ensembles; HE: Hacking Metrics; RAI: Responsible AI; PM: Predictive Multiplicity; EM: Explanation Multiplicity; OM: Other Multiplicity; RSE: Rashomon Set Exploration; REML: Rashomon Effect in ML; App: Application; Sur: Survey.
2. The Rashomon Effect in Machine Learning
Taking its name from Akira Kurosawa’s 1950 film Rashomon, the Rashomon effect is an epistemological framework that highlights the subjectivity and ambiguity inherent in human perception (Anderson, 2016; Davis et al., 2015; Alia, 1996; Heider, 1988). Borrowing from Davis et al. (2015), the Rashomon effect can be defined as “a combination of a difference of perspective and equally plausible accounts, with the absence of evidence to elevate one above others, […]”. The Rashomon effect has been studied in several different domains, like the influence of cognitive biases on memory (Tindale, 2016; Trabasso, 2018), the impact of culture and the fluidity of truth in ethnographic studies (Heider, 1988), the study of context, medium, and framing of communication (Anderson, 2016; Soreanu and German, 2022), the unreliability of eyewitnesses (Pansky et al., 2005; Hirst and Brown, 2011), and–central to our discussion–algorithmic modelling and machine learning (Breiman, 2001; Black et al., 2022).
The term Rashomon effect was first introduced into algorithmic modelling by Breiman (2001), pointing out the presence of a set of good models that all achieve similar error rates. It has since been used in discussions of statistical modelling (Wang and Tao, 2008; Bonate, 2006; Broadhurst and Kell, 2006; Ueki and Kawasaki, 2013), null hacking (Protzko, 2018), designing robust algorithms (Tulabandhula and Rudin, 2014; Castillo et al., 2008), measuring variable importance (Dong and Rudin, 2019; Fisher et al., 2019), and applications in various domains (Kang et al., 2018; Chantre et al., 2018; Sreedharan et al., 2018). More recently, it has found a resurgence with increasing attention given to multiplicity in machine learning, evident both in studies that directly address the topic (Marx et al., 2020; Black et al., 2022; D’Amour et al., 2022; Rudin et al., 2024; Paes et al., 2023; Del Giudice, 2024; Biecek et al., 2024) and in research that situates multiplicity within the broader context of other fields (Møllersen and Holsbø, 2023; Rudin et al., 2022; Molnar et al., 2020; Jiang et al., 2024b; Biecek and Samek, 2024).
Why do we see the Rashomon effect? In machine learning, data serves as a proxy for the real world, yet it inherently loses information at multiple stages. The first step—translating the world into a data generation process—simplifies complex relationships, introducing randomness to account for uncontrollable aspects. Zhang et al. (2020) termed this “distributional complexity” (associated with ‘aleatoric’–Latin aleatorius–meaning “dice” or “game of chance”), reflecting the challenge of how well the distribution represents the real world. However, even this distribution remains out of our reach; instead, we work with finite samples. This second step of information loss, described as ”approximation complexity” (associated with ‘epistemic’–Latin episteme–meaning “knowledge”) (Zhang et al., 2020), relates to how well can we approximate the underlying generation using finite data. Unlike distributional complexity, which is irreducible, approximation complexity can be mitigated through better data quality and improved algorithms.
Together, distributional and approximation complexities define the fundamental loss of information in machine learning, resulting in gaps where multiple interpretations—and thus the Rashomon effect—can arise. While viewing the Rashomon effect through this lens of information loss is insightful, it does not provide an operational framework for developers to examine or address these challenges. Therefore, in this section, we instead focus on the role of developer choices in model design, laying the foundation for discussions on multiplicity.
Figure 2. Impact of developer choices on individuals and the population downstream.
2.1. Rashomon Effect through Design Choices and Model Selection
Designing a machine learning model involves a series of interconnected or choices. Beginning with the data, decisions are made regarding how to process and filter data, which features to select, etc. (Meyer et al., 2023; Simson et al., 2024; Cavus and Biecek, 2024b) Even the random partitioning of the data into training and validation impacts final model behavior (Friedler et al., 2019). Beyond data, the learning algorithm design further entails numerous decisions: model architecture (Arnold et al., 2024; Rudin et al., 2024), hyperparameters (Bouthillier et al., 2021; Arnold et al., 2024), various forms of stochasticity (Picard, 2021; Bouthillier et al., 2021; Pecher et al., 2024; Sellam et al., [n. d.]; McCoy et al., 2020), and even the evaluation and model selection criteria (Ganesh, 2024; Ganesh et al., 2024). Each decision contributes to the cascade of choices that directly impacts the multiplicity of the trained models (see Figure 2).
Notably, these choices are not always well-informed. In some cases, they are intentional, guided by insights from the literature on the effects of algorithm design on model behaviour (Wu et al., 2021; Ponomareva et al., 2023; Kulynych et al., 2023; Ganesh, 2024). In others, they are conventional, driven by popular trends or convenience (Shwartz-Ziv and Armon, 2022; Dubey et al., 2022; Creel and Hellman, 2022; Sellam et al., [n. d.]). Finally, some choices remain arbitrary, like choosing a random seed. It is through training multiple models and evaluating them, that we grasp the impact of these arbitrary choices. To connect these choices with key subdomains in the multiplicity literature, we introduce the Intent-Convention-Arbitrariness (ICA) framework. More specifically, we argue that these choices exhibit the following properties:
  • Intentional Choices and Steering Model Behaviour: We define certain choices as intentional when developers make deliberate choices with an understanding of their impact on the final model to achieve desired outcomes. Examples include incorporating bias-mitigating regularization to enhance fairness (Hort et al., 2024; Chen et al., 2023), using simpler models for better interpretability (Rudin et al., 2024), or applying data augmentation to improve robustness (Rebuffi et al., 2021; Eghbal-zadeh et al., 2024). These choices are made to steer model behaviour in a particular direction, giving the developer control over navigating the Rashomon set without the need to train multiple models, i.e., avoiding model selection. Intentional choices are typically informed by extensive prior research or other advancements that guide the developer. Note that choices like selecting a pre-trained model because it is the only available option do not qualify as an intentional choice. Although the developer is aware of the impact of their choice, the decision is made out of necessity rather than deliberate intent.
  • Conventional Choices and Homogenization: We define conventional choices as choices made without knowledge of their impact, out of convenience, or due to lack of alternatives. Examples include adopting popular models or hyperparameters without evaluating their suitability for the specific application (Shwartz-Ziv and Armon, 2022; Dubey et al., 2022), such as using neural networks where simpler models would suffice (Rudin, 2019; Shwartz-Ziv and Armon, 2022), or applying out-of-the-box fairness, robustness, or explanation techniques without understanding their implications (Gichoya et al., 2023; Rudin, 2019; Adebayo et al., 2018; Lipton, 2018; Tramer et al., 2020; Carlini et al., 2019). By definition, conventional choices follow established norms or trends within the field rather than addressing specific needs. As a result, these choices contribute to “homogenization”, where models trained by different developers exhibit similar behaviour, and can introduce systemic harm (Creel and Hellman, 2022; Wu et al., 2024; Kleinberg and Raghavan, 2021; Jain et al., 2024b; Bommasani et al., 2022). As shown by Bommasani et al. (2022), even shared components, for instance, common choices across developers, can lead to homogenized outcomes across multiple systems (§6.2).
  • Arbitrary Choices and Model Selection: Unlike intentional or conventional choices, arbitrary choices have an indeterminate relationship with the final model, usually evaluated after training the model. Examples include choosing the random seed for weight initialization, data shuffling, dropout, etc., or arbitrary hyperparameter variations such as selecting between a learning rate of 0.008, 0.01, or 0.012. When dealing with arbitrary choices, making appropriate choices depends on model selection post-training. This has been widely recommended in the literature for auditing multiplicity, creating ensembles, or navigating among competing choices (Kulynych et al., 2023; Long et al., 2024; Creel and Hellman, 2022). However, training multiple models can be expensive, especially for large models or complex hypothesis classes (Hsu et al., 2024, [n. d.]; Kissel and Mentch, 2024; Zhong et al., 2024; Mata et al., 2022; Ciaperoni et al., 2024). Even when feasible, model selection risks overfitting, potentially undermining generalization (Ganesh, 2024; Cooper et al., 2021; Coker et al., 2021). These challenges don’t necessarily dismiss the efficacy of training multiple models. Rather, they highlight the complexities of this approach and motivate a deeper investigation into the benefits and pitfalls of model selection (§5).
In practice, few choices, if any, fall entirely into a single category, Instead, every choice made by a developer involves a mix of intentional, conventional, and arbitrary factors. For instance, consider a developer using dropout testing dropout parameter values 0.1, 0.2, and 0.3. The decision to use dropout is likely intentional, informed by a goal to enhance model robustness. However, the choice of dropout over other regularization techniques may be conventional, driven by its widespread use or integration in popular frameworks. Finally, the specific dropout rates chosen are arbitrary, as their impact on the model’s performance is determined through selection post-training. Understanding the balance between these three factors is essential for navigating the challenges posed by the Rashomon effect.
2.2. Impact of Rashomon Effect
With a grasp of how the Rashomon effect manifests in machine learning, we turn to its impact, i.e., we introduce multiplicity. While the existence of multiple good models is undeniably intriguing, the Rashomon sets should not be a mere academic curiosity. These sets only matter in context, where a change of perspective by choosing a different model from the set, i.e., a different developer choice, influences real-world outcomes (see example in Figure 2). Multiplicity is, thus, the variation in model behaviour across the Rashomon set that holds contextual value. Here, we provide a brief intuition of the various forms of multiplicity, setting the stage for its formal definition in the next section.
The most straightforward example of multiplicity is when individuals encounter conflicting predictions from models in the Rashomon set, known as predictive multiplicity (Marx et al., 2020). Such conflicts create brittleness and arbitrariness in decision-making, undermining the reliability of these models and hampering effective planning (D’Amour et al., 2022; Marx et al., 2020; Black et al., 2022; Cooper et al., 2022; Watson-Daniels et al., 2024; Milani Fard et al., 2016). While predictive multiplicity can be harmful in critical domains, such as medical or legal decisions, it is, however, not inherently bad. For instance, purposefully controlled arbitrariness (called ‘randomness’ to distinguish from uncontrolled arbitrariness) in predictions can help address the concerns of outcome homogenization (Jain et al., 2024b, a; Creel and Hellman, 2022; Barocas et al., 2023; Perry and Zarsky, 2015).
Unsurprisingly, predictive multiplicity has received significant attention in the literature ($35\%$ of papers in our systematic review; see Figure 1). Yet, this is only one aspect of the Rashomon effect. Consider, for instance, the rapidly growing field of inconsistency in explanations provided by models within the Rashomon set ($37.5\%$ of papers in our systematic review; see Figure 1). Studies have shown that models in the Rashomon set often produce conflicting variable importance and feature attribution scores (Black et al., 2021; Laberge et al., 2023; Hancox-Li, 2020; Baniecki et al., 2024; Li et al., 2024; Gunasekaran et al., 2024; Müller et al., 2023; Poiret et al., 2023; Okazaki et al., 2024; Tan, 2024; Kuzilek and Çavuş, 2024; Kumar et al., 2023), which can undermine trust, for example, confusing clinicians that rely on explanations during AI-assisted diagnostics. Similarly, counterfactuals generated by one model in the Rashomon set often fail to transfer to others (Pawelczyk et al., 2020; Leofante et al., 2023; Jiang et al., 2024a; Hasan and Talbert, 2022). This poses significant challenges for algorithmic recourse when models are regularly updated, as recourse provided by one model may become invalid when replaced by another, questioning their legitimacy and undermining user trust (Rawal et al., 2021; Leofante et al., 2023).
More broadly, the effects of the Rashomon effect—multiplicity—can be measured as any form of behavioural difference between models in the context of real-world impact. These effects are multifaceted, and while predictive and explanation multiplicities are the most recognized, narrowing our focus underestimates the broader risks associated with other underrepresented forms of multiplicity (only $\sim 24\%$ of papers in our systematic review cover other forms of multiplicity; see Figure 1). In the next section, we will formally define both Rashomon sets and multiplicity, expanding upon existing definitions in the literature and drawing inspiration from the discussion presented here.
3. Definitions and Metrics
We now define Rashomon sets and multiplicity, building on existing literature while expanding the scope to encompass a wider range of works. Additionally, we review multiplicity metrics and the literature on evaluating multiplicity.
3.1. Formalizing Rashomon Sets and Multiplicity
The concept of multiplicity is deeply rooted in the Rashomon effect, and the models illustrating this Rashomon effect are together known as a Rashomon set, a set of competing models, a set of good models, $\epsilon$-Rashomon set, $\epsilon$-Level set, etc. We’ll stick with the term Rashomon set for consistency in our formalization. In essence, the Rashomon set represents a set of models that are practically indistinguishable, underscoring the arbitrariness in choosing one model over another. Thus, we need to begin by defining these models and explaining what it means for them to be indistinguishable.
Existing work in multiplicity tends to adopt a narrow view of these models. For instance, much of the research that formalizes the multiplicity problem restricts model choices to a specific hypothesis class $\mathcal{H}$ and/or limits them to training on a fixed, pre-processed dataset $\mathcal{D}$ (Marx et al., 2020; Castillo et al., 2008; Teney et al., 2022; Watson-Daniels et al., 2023b, a). However, this overlooks developer choices made during data collection, data processing, and even the model training stages, all of which can influence the final model, as discussed above. Similarly, most existing studies define indistinguishability solely in terms of loss (Castillo et al., 2008; Teney et al., 2022; Paes et al., 2023; Du et al., 2024; Hamman et al., 2024), disregarding other crucial choices involved in designing evaluation criteria and model selection, as previously mentioned.
To explicitly broaden the definition of Rashomon sets, we introduce a set of metric delta functions, $\mathbf{\Delta}^{P}$, and corresponding thresholds $\mathcal{E}^{P}$. A metric delta function takes as input two models and measures the difference between them under the given metric. These metric deltas determine whether two models are indistinguishable: if the difference in performance for every $\delta^{P}_{i}\in\mathbf{\Delta}^{P}$ falls within the corresponding threshold $\epsilon^{P}_{i}\in\mathcal{E}^{P}$. For example, one might define the metric delta for accuracy as $\delta_{Acc,D}(h_{1},h_{2})=|Acc(h_{1},D)-Acc(h_{2},D)|$. Formally,
Definition 3.1 (Rashomon Set).
Two models $h_{1},h_{2}$ belong to the same Rashomon set under performance constraints $(\mathbf{\Delta}^{P},\mathcal{E}^{P})$ if they exhibit similar performance for every metric in the given performance constraints, i.e.:
(1) $\delta^{P}_{i}(h_{1},h_{2})\leq\epsilon^{P}_{i}\;\;\;\;\forall\;(\delta^{P}_{i},\epsilon^{P}_{i})\in(\mathbf{\Delta}^{P},\mathcal{E}^{P})$
We now turn to defining the context in which these models exhibit diverse behaviour, i.e., multiplicity. Again, as previously noted, most existing work focuses on either predictive or explanation multiplicity (see Figure 1), with limited attention given to other forms such as fairness multiplicity (Islam et al., 2021; Coston et al., 2021; Ganesh et al., 2023; Gillis et al., 2024; Sellam et al., [n. d.]; Wang and Russakovsky, 2021), OOD robustness multiplicity (Teney et al., 2022; McCoy et al., 2020), model complexity multiplicity (Semenova et al., 2024; Liu et al., 2022; Rudin et al., 2024; Semenova et al., 2022; Coupkova and Boutin, 2024), and feature interaction multiplicity (Li et al., [n. d.]), among others. To capture the various impacts discussed above, we generalize multiplicity by binding it to a metric delta $\delta^{M}$ and corresponding threshold $\epsilon^{M}$, on models belonging to the Rashomon set. Formally,
Definition 3.2 (Model Multiplicity).
Two models $h_{1},h_{2}$ exhibit multiplicity under performance constraints $(\mathbf{\Delta}^{P},\mathcal{E}^{P})$ and multiplicity constraint $(\delta^{M},\epsilon^{M})$, if they have similar performance for every metric in the performance constraints yet differ on the metric in the multiplicity constraint, i.e.:
(2) $\delta^{P}_{i}(h_{1},h_{2})\leq\epsilon^{P}_{i}\;\;\;\;\forall\;(\delta^{P}_{i},\epsilon^{P}_{i})\in(\mathbf{\Delta}^{P},\mathcal{E}^{P})\quad\quad\textrm{and}\quad\delta^{M}(h_{1},h_{2})>\epsilon^{M}\end{aligned}$
Although our definitions are quite similar to those in existing literature, we have deliberately generalized them to encompass a broader range of underrepresented works. These are not radical changes, but we believe are crucial in drawing attention to various overlooked choices in the model design pipeline and to better understand their role in multiplicity. However, we also recognize that in certain contexts, a specific definition of multiplicity might be needed. In such contexts, our definition can be reduced appropriately to match the use case.
3.2. Evaluating Multiplicity
Metric Original Objective Problem Setting Monotonic Resolution
Ambiguity (Marx et al., 2020) Predictive Multiplicity Multi-Class Classification Yes Dataset
Obscurity (Cavus and Biecek, 2024a) Predictive Multiplicity Multi-Class Classification No Dataset
Discrepancy (Marx et al., 2020) Predictive Multiplicity Multi-Class Classification Yes Dataset
Indistinguishability (Paes et al., 2023) Predictive Multiplicity Multi-Class Classification - Dataset
Degree of Underspecification (Teney et al., 2022) Predictive Multiplicity Multi-Class Classification Yes Dataset
Viable Prediction Range (Watson-Daniels et al., 2023b) Predictive Multiplicity Probabilistic Classification Yes Individual
Rashomon Capacity (Hsu and Calmon, 2022) Predictive Multiplicity Probabilistic Classification Yes Individual
Multi-target Ambiguity (Watson-Daniels et al., 2023a) Predictive Multiplicity Multi-Target Classification Yes Dataset
Rank List Sensitivity (Oh et al., 2022) Predictive Multiplicity Recommender Systems - Dataset
Std. of Scores (Heljakka et al., 2022; Long et al., 2024) Predictive Multiplicity Agnostic No Dataset
Representational Multiplicity (Heljakka et al., 2022) Procedural Multiplicity Agnostic No Dataset
Region Similarity Score (Somepalli et al., 2022) Procedural Multiplicity Agnostic - Dataset
Variance Arbitrariness Agnostic No Individual
Self-consistency (Cooper et al., 2024) Arbitrariness Multi-Class Classification No Individual
$\epsilon$-robust to Dataset Multiplicity (Meyer et al., 2023) Dataset Multiplicity Regression - Dataset
Unfairness Range (Mata et al., 2022) Fairness Multiplicity Agnostic Yes Dataset
Rashomon Ratio (Semenova et al., 2022) Size of Rashomon Set Agnostic Yes Dataset
Underspecification Score (Madras et al., 2019) Underspecification Multi-Class Classification No Individual
Accuracy Under Intervention (Ganesh, 2024) Metric Multiplicity Multi-Class Classification - Dataset
Metric Original Objective Explanation Technique Monotonic Resolution
Consistency (Watson et al., 2022) Explanation Multiplicity Agnostic No Dataset
Model Class Reliance (Fisher et al., 2019) Explanation Multiplicity Model Reliance (Fisher et al., 2019) Yes Dataset
Attribution Agreement (Krishna et al., 2022; Brunet et al., 2022) Explanation Multiplicity Feature Attribution - Individual
Profile Disparity Index (Kobylińska et al., 2024) Explanation Multiplicity Model Profile (Kobylińska et al., 2024) - Dataset
Inv. Cost of Neg. Surprise (Pawelczyk et al., 2020) Explanation Multiplicity Counterfactuals - Dataset
Variable Importance Clouds (Dong and Rudin, 2019) Explanation Multiplicity Feature Attribution Yes Dataset
Coverage & Interval Width  (Marx et al., 2023) Explanation Uncertainty Agnostic No Individual
Table 1. Existing multiplicity metrics in the literature, along with original objective, problem setting, monotonicity wrt to the Rashomon set, and metric resolution. Metrics with no entry under ‘Monotonic’ are defined to compare two models, and not the Rashomon set.
To complement the expanded definitions, we compile a comprehensive list of metrics from the literature that measure various forms of multiplicity in Table 1. For each metric, we mark its original objective, the problem setting, resolution (i.e., whether the metric applies to individual data points or the entire dataset), and if it exhibits set monotonicity within the Rashomon set (i.e., whether a reduction in the Rashomon set size implies a monotonic change in the metric). Monotonicity in a metric is particularly desirable in certain scenarios (Hsu and Calmon, 2022), because it ensures that reducing the Rashomon set size—by improving specification—will either decrease or maintain the multiplicity. However, in other contexts, metrics that are not necessarily monotonic, such as variance, can be equally valuable. For instance, when it is important to quantify not just the range but also the dispersion of arbitrariness, metrics that capture such variability become crucial. Ultimately, the relevance of monotonicity in a metric depends on the specific context in which it is applied. By systematically recording different facets of each metric in Table 1, we hope to provide a structure for practitioners to identify the most suitable metric for their specific needs in evaluating multiplicity.
Beyond the metrics themselves, we also revisit the challenge of training multiple models to evaluate multiplicity, exploring more efficient alternatives. Madras et al. (2019) propose a local ensembling technique to quantify underspecification by analyzing the loss curvature, eliminating the need to train multiple models. Hsu et al. (2024) use a similar local approach and show that Monte Carlo dropout can be adopted to approximate multiplicity when constrained by utility considerations (i.e., Rashomon sets). In contrast to these local methods, Kissel and Mentch (2024) introduce a model path selection technique that incrementally builds from simpler to more complex models. This approach efficiently constructs the Rashomon set by recursively increasing the complexity of plausible models. Other model class-specific techniques have also been proposed to explore Rashomon sets more efficiently (Hsu et al., [n. d.]; Mata et al., 2022; Zhong et al., 2024; Xin et al., 2022). Despite these advancements, much work remains to be done to make the enumeration of Rashomon sets more efficient and practical.
4. Multiplicity, Uncertainty and Bias-Variance Decomposition
When examining arbitrariness in decision-making, machine learning research often focuses on prediction uncertainty (Gal, 2016; Gal et al., 2017; Kendall and Gal, 2017; Smith et al., [n. d.])–model’s lack of confidence in its predictions–or the bias-variance decomposition (Domingos, 2000; Kong and Dietterich, 1995; BREIMAN, 1996)–dividing the model error into how well the model fits the data (bias) and its sensitivity to changes in data (variance). With extensive literature already present in these areas, a natural question arises: What unique perspectives does multiplicity bring to the discussion of arbitrariness not already covered by these concepts? In this section, we formalize the interplay between multiplicity, uncertainty, and bias-variance decomposition, addressing this question mathematically and through practical recommendations for when different perspectives on arbitrariness are most valuable.
4.1. Multiplicity and Uncertainty
Prediction Uncertainty: We start by defining uncertainty, drawing heavily from Gal (2016); Gal et al. (2017); Kendall and Gal (2017); Smith et al. ([n. d.]). Prediction uncertainty quantifies the degree of confidence—or lack thereof—in a model’s predictions. As it reflects the lack of confidence in a model’s predictions, uncertainty is often represented as the randomness (or entropy) in those predictions. Formally, prediction uncertainty is commonly defined as:
(3) $U(x,D)={\textrm{H}}_{y}[{\textrm{Prob}}(y|x,D)]$ $={\textrm{H}}_{y}[\Sigma_{\theta}[{\textrm{Prob}}(y|x,\theta)*{\textrm{Prob}}(\theta|D)]]$
(4) $={\textrm{H}}_{y}[{\textrm{E}}_{\theta\sim{\textrm{Prob}}(\theta|D)}[{\textrm{Prob}}(y|x,\theta)]]$
[…]
Prediction uncertainty is also typically divided into aleatoric and epistemic uncertainty (Hüllermeier and Waegeman, 2021; Senge et al., 2014; Kendall and Gal, 2017). Aleatoric uncertainty, connected to “distributional complexity” (Zhang et al., 2020), is the intrinsic difficulty of modelling the relationship in the data distribution, often described as the uncertainty left with access to infinite data. In contrast, epistemic uncertainty, connected to “approximation complexity” (Zhang et al., 2020), stems from a lack of knowledge about the best model to use, i.e., the challenge of accurately capturing the true data distribution. Epistemic uncertainty does not have a straightforward formulation. Instead, it is commonly defined as the uncertainty remaining after accounting for aleatoric uncertainty.
For the interested reader, we encourage exploring the uncertainty literature (Hüllermeier and Waegeman, 2021; Tran et al., 2020; Kendall and Gal, 2017; Gal, 2016), as we do not expand on these concepts in our work. We simply restate these definitions to compare them directly with multiplicity.
Predictive Multiplicity through the lens of Uncertainty: We temporarily redefine multiplicity, drawing on the same principles used to define uncertainty. In simple terms, we also define multiplicity as the entropy of predictions, but only limited to models within the Rashomon set $R$. Thus, we can formalize multiplicity $M(x,D)$ as:
(7) $M(x,D)={\textrm{H}}_{y}[{\textrm{E}}_{\theta\sim{\textrm{Prob}}_{R}(\theta|D)}[{\textrm{Prob}}(y|x,\theta)]]\quad\quad\textrm{s.t.}\quad{\textrm{Prob}}_{R}(\theta|D)=\left\{\begin{array}[]{lr}0&\textrm{if}\;\theta\notin R\\ \dfrac{{\textrm{Prob}}(\theta|D)}{\sum_{\theta\in R}{\textrm{Prob}}(\theta|D)}&\textrm{if}\;\theta\in R\end{array}\right\}$
where ${\textrm{Prob}}_{R}(\theta|D)$ represents a modified probability distribution that only includes the models in the Rashomon set $R$.
Comparing equations 4 and 7, it is clear that the expectation terms are defined over different distributions: over all possible models for uncertainty (eq 4), and over only models within the Rashomon set for multiplicity (eq 7). But why does this distinction matter, and why should we care about both? To answer this, we discuss practical scenarios where viewing a problem through the lens of multiplicity is more appropriate than uncertainty and vice-versa.
Uncertainty or Multiplicity? Choosing the Right Lens. Multiplicity examines prediction consistency, while uncertainty assesses confidence. Uncertainty is better suited for modelling the information-theoretic relationship between data and the predictions derived from it. Multiplicity, on the other hand, offers the most relevant perspective for actively exploring the various interpretations that can emerge during learning. Similarly, when examining how different modelling choices or model selection criteria can influence outcomes, the lens of multiplicity proves invaluable.
We outline some characteristics to look for when deciding whether to use the lens of multiplicity or uncertainty.
  • Uncertainty provides an information-centric perspective. As uncertainty definitions are derived from information theory (Box and Tiao, 1973; Kendall and Gal, 2017; Gal, 2016), it is a fundamentally better fit for related analyses. For example, uncertainty plays a crucial role in active learning, by finding instances most likely to provide maximum new ‘information’ (Yang and Loog, 2016; Sharma and Bilgic, 2017; Nguyen et al., 2022).
  • Uncertainty is sufficient when dealing with distributional complexity. Noise in real-world data may result in a lack of predictive power to make reliable decisions (Wang et al., 2024). Having access to different interpretations through multiplicity adds little value in such cases—multiple ways of delivering incorrect answers do not necessarily enhance usefulness.
  • Uncertainty quantification can be more efficient, but research in multiplicity quantification is growing rapidly. Uncertainty is streamlined into modern machine learning pipelines through Bayesian networks and model calibration (Jospin et al., 2022; Niculescu-Mizil and Caruana, 2005; Vaicenavicius et al., 2019), offering a cost-effective alternative to multiplicity. Even when training multiple models, there is typically no definitive way to ensure that every trained model falls within the Rashomon set, and thus not all trained models contribute to measuring multiplicity whereas they are still valuable for quantifying uncertainty. That said, advancements in multiplicity research, as previously discussed in §3.2, have already been challenging this dynamic and may continue to reshape it further in the future (Madras et al., 2019; Hsu et al., 2024; Kissel and Mentch, 2024; Hsu et al., [n. d.]; Mata et al., 2022; Zhong et al., 2024).
  • Multiplicity is aligned with learning theory and hierarchical optimization. Every decision in the learning algorithm influences the underlying optimization. Multiplicity can help scrutinize how each choice shapes the final model. Applications include the impact of data processing, random seeds, hyperparameters, etc. (Islam et al., 2021; Ganesh et al., 2023; Meyer et al., 2023; Ganesh, 2024; Cavus and Biecek, 2024b), or, broadly, any form of bi-level or constrained optimization, including meta-objectives (Black et al., 2024b; Semenova et al., 2024; Sun et al., [n. d.]; Ganesh et al., 2023; Gerchick et al., 2023).
  • Multiplicity is better suited to explore alternative interpretations. As we’ve discussed, choosing among different learned interpretations can introduce arbitrariness. Multiplicity, particularly Rashomon sets, enables exploration of these alternative interpretations. Examples include personalization with model mixtures, combining multiple models, homogenization concerns, etc. (Ma et al., 2020; Creel and Hellman, 2022; Wu et al., 2024; Bommasani et al., 2022; Black et al., 2021; Laberge et al., 2023; Breiman, 2001; Long et al., 2024)
Note that our recommendations paint a broad picture of when the lens of multiplicity or uncertainty could be useful, but these are intended only as guidelines and deviation from these may be warranted in specific contexts.
4.2. Multiplicity and Bias-Variance Decomposition
Another common measure of arbitrariness in decisions is the ‘variance’ from the bias-variance decomposition. Error in machine learning is often categorized into three parts: irreducible error, bias, and variance (Domingos, 2000; Kong and Dietterich, 1995; BREIMAN, 1996). This is known as the bias-variance decomposition. The terms ‘bias’ and ‘variance’ describe how well the model approximates the underlying distribution, i.e., approximation complexity, where ‘bias’ captures how well the chosen model fits the given data, while ‘variance’ reflects the model’s sensitivity to variations in the dataset.
Understandably, at first glance, sensitivity to variations in the dataset—captured by ‘variance’—might seem like a natural way to measure arbitrariness in decision-making. However, bias-variance decomposition is typically confined to the analysis of a single model, focusing only on sensitivity to variations in the underlying dataset. While this framework provides valuable insights into a model’s behaviour—guiding preferences for models with lower bias or variance—it does not address the broader arbitrariness introduced by the entire learning pipeline. Multiplicity, in contrast, enables comparisons across different models generated through the pipeline. Formally,
(8) $Error=IrreducibleError+Bias+{}$ $\mathbf{Var_{D}[f_{\theta}(x)]}$  (Domingos, 2000; Kong and Dietterich, 1995; BREIMAN, 1996)
(9) $PredictiveMultiplicity={}$ $\mathbf{Var_{\theta\sim{\textrm{Prob}}_{R}(\theta|D)}[f_{\theta}(x)]}$  (Heljakka et al., 2022; Long et al., 2024; Hamman et al., 2024)
where $f_{\theta}(.)$ is the predictive function for the learned parameter value $\theta$, and $Var[.]$ represents variance. We use a metric of predictive multiplicity that also uses variance to quantify multiplicity, facilitating clearer comparisons (Heljakka et al., 2022; Long et al., 2024; Hamman et al., 2024). While the concept of ‘variance’ from bias-variance decomposition focuses on a single model’s sensitivity to variations in the dataset, multiplicity, in contrast, captures changing predictions across multiple models. As we will explore further in §5, multiplicity allows us to navigate through multiple candidate models and even identify those with desirable bias-variance properties–such as ‘simpler’ models, often associated with lower variance (Semenova et al., 2024; Liu et al., 2022; Rudin et al., 2024; Semenova et al., 2022; Coupkova and Boutin, 2024).
We showed that multiplicity is neither universally redundant nor superior when compared to uncertainty or bias-variance decomposition. Instead, we offer guidance on when each approach is most appropriate. It is also crucial to recognize the common approximations made in the literature and to explicitly acknowledge their assumptions. For instance, predictive multiplicity has been used in the literature to measure epidemic uncertainty (Ali et al., 2021; Shahbab and Zaman, 2024; Brunet et al., 2022). This stems from the intuition that epistemic uncertainty is often estimated using Bayesian neural networks, whose behaviour can resemble Rashomon sets. However, this approximation is only valid if the models are calibrated, meaning they already account for aleatoric uncertainty (Hüllermeier and Waegeman, 2021). Furthermore, it is important to distinguish between multiplicity within the Rashomon set and uncertainty (of any kind) across the entire set of models, as discussed earlier. Thus, using multiplicity as a proxy for epistemic uncertainty is an approximation which may fail in situations with high aleatoric uncertainty or when the distribution over the Rashomon set is significantly different from that on the entire set of models.
The interplay between multiplicity, uncertainty, and bias-variance decomposition remains complex, underscoring the need for further research to better understand and effectively utilize these concepts.
5. Exploring Alternative Interpretations with Multiplicity
Now that we’ve discussed the formalization of multiplicity, let’s delve into its real-world implications. One of the biggest advantages of multiplicity lies in its ability to uncover and explore various ‘good’ learned interpretations (Black et al., 2022; Rudin et al., 2024). When multiple interpretations exist, it is reasonable to expect that some of them may exhibit certain desirable properties, such as better fairness, robustness, interpretability, etc. Additionally, moving beyond a single-model paradigm also opens the door to aggregating insights from multiple models. In this section, we study how multiplicity facilitates such exploration and its broader implications. Our discussion builds on what Black et al. (2022) refer to as the aggregate-level effects of multiplicity, while incorporating insights from more recent developments in the field.
5.1. Searching Instead of Optimizing
Machine learning often involves tackling complex optimizations, including two commonly seen hierarchical optimization problems: bi-level optimization and constrained optimization. Bi-level optimization refers to scenarios where one optimization problem depends on variables governed by another nested optimization (Sinha et al., 2017; Zhang et al., 2024). A classic example is hyperparameter optimization. Constrained optimization, on the other hand, involves solving the optimization problems under specific constraints (Bertsekas, 2014; Gallego-Posada, 2024; Goodfellow et al., 2014). Examples include enforcing constraints of fairness, robustness, safety, etc. during model training (Gallego-Posada, 2024; Dwork et al., 2012). These classes of optimization problems can be notoriously difficult to solve. Challenges arise from the complexity of formalizing certain constraints, the difficulty of creating continuous and differentiable relaxations, the absence of closed-form solutions, or evolving client requirements, among others.
Interestingly, multiplicity offers a practical workaround: if you can’t optimize, search for it! Brute-force strategies of searching through various potential solutions are well-established in both bi-level optimization and constrained optimization literature (Sinha et al., 2017; Zhang et al., 2024; Gallego-Posada, 2024). Though more efficient reductions and mathematical alternatives are preferred when feasible, they are not always possible. In such cases, searching through potential solutions to find the best fit becomes a viable strategy, and multiplicity plays a pivotal role in this search (see Figure 3). Multiplicity has been used in the literature to find models with lower bias (Black et al., 2024b; Simson et al., 2024; Ganesh et al., 2024; Islam et al., 2021), smaller model complexity (Semenova et al., 2022, 2024; Boner et al., [n. d.]; Rudin et al., 2024; Doctor et al., 2024; Coupkova and Boutin, 2024), better explanations (Sun et al., [n. d.]; Veran et al., 2023), improved generalizability (Li et al., 2024), and the ability to allow personalization (Ma et al., 2020). Beyond simply enumerating the Rashomon set to search for better models, several recent works have also shown how the visualization of the Rashomon set can empower developers to select models that meet their specific requirements (He and Shaposhnik, 2023; Eerlings et al., 2024; Luyten et al., 2024).
Figure 3. The role of multiplicity in brute-force search for bi-level and constrained optimization problems in machine learning.
5.2. Ensembles and More
Selecting the “best” model may not always be recommended, particularly when no single interpretation of the data can be optimal. Instead, combining insights from multiple models can often yield better results in such scenarios. Techniques like prediction ensembling, or bagging, have long been a central recommendation for stability in machine learning (Breiman, 1996; Dietterich, 2000; Long et al., 2024). Many methods in the multiplicity literature have capitalized on combining various forms of information from models in the Rashomon set. While literature in this direction primarily focuses on aggregating model explanations to create more stable and reliable explanations (Marx et al., 2023; Black et al., 2021; Jiang et al., 2024a; Laberge et al., 2023; Ley et al., 2023; Sreedharan et al., 2018; Baniecki et al., 2024; Gunasekaran et al., 2024; Mitruț et al., 2024; Hibbeln et al., 2024; Donnelly et al., 2023; Fisher et al., 2019; Dong and Rudin, 2020; Kobylińska et al., 2024; Zuin et al., 2021; Hamamoto and Egi, 2021; Kowal, 2022; Li et al., [n. d.]; Smith et al., 2020), other works have also shown the benefits of aggregating fairness scores (Coston et al., 2021), individual probabilities (Roth et al., 2023), or regression analysis to discover causality (Ueki and Kawasaki, 2013). These techniques demonstrate the value of leveraging multiplicity not just to select a single best model but instead to combine multiple learned interpretations.
5.3. Hacking Metrics with Multiplicity
While multiplicity can enable the discovery of better models, it also introduces risks, particularly the potential of exploiting these search methods to circumvent regulatory requirements and interventions. This is more prevalent when broad principles, such as fairness, are reduced to specific benchmarks or metrics (Black et al., 2024a). By leveraging multiplicity, it becomes possible to ‘hack’ these metrics, producing models that meet the specified criteria without truly adhering to the underlying principles the metrics are intended to enforce (Coker et al., 2021; Cooper et al., 2021; Møllersen and Holsbø, 2023). Several studies in the literature have shown that such a search can indeed produce models capable of “regulatory-washing”, being able to manipulate explanations (Sharma et al., 2024; Aïvodji et al., 2021; Shahin Shamsabadi et al., 2022) and fairness scores (Forde et al., 2021; Black et al., 2024a; Ganesh et al., 2024). Such manipulation can also occur unintentionally-a result of overfitting to a given metric (Black et al., 2024a; Cooper et al., 2021; Ganesh, 2024)-underscoring the need for vigilance against the misuse of multiplicity and calling for a more robust operationalization of regulatory frameworks (Black et al., 2024a).
6. Multiplicity and Responsible AI
We now move to placing multiplicity in the broader landscape of responsible AI. In this section, we examine the two key concerns for individuals originating from multiplicity, i.e., arbitrariness in model selection and outcome homogenization. Again, our discussion here builds on what was termed as the individual-level effects of multiplicity by Black et al. (2022), while also focusing on topics that have gained more interest recently, such as homogenization.
6.1. Arbitrariness as a Responsible AI Concern
Arbitrary decisions in an automated system can be deeply concerning, particularly in critical domains such as law and medicine, where they can have direct and lasting impacts on human lives (Black et al., 2022; Gomez et al., 2024; Sokol et al., 2024; Barocas et al., 2023; D’Amour et al., 2022; Watson-Daniels et al., 2024). Borrowing an analogy from Gomez et al. (2024), imagine a judge deciding legal cases by flipping a coin. While this may seem extreme, it demonstrates how machine learning models can have arbitrariness embedded in them due to an analogous coin flip done by the developer during model design. This aligns closely with our ICA framework (§2.1), where we discuss the arbitrary choices by the developers that can contribute to multiplicity.
There are contexts where a degree of “controlled randomness” may be acceptable—or even necessary (§6.2). However, arbitrariness is a significant concern in scenarios where individuals lack access to other ‘equivalent opportunities’ (Barocas et al., 2023). For instance, in hiring, some level of arbitrariness may be acceptable, or even useful to deal with the concerns of homogenization (§6.2). This is because individuals looking for a job often seek multiple opportunities across different companies, increasing their chances of being hired elsewhere. In contrast, domains such as law or medicine typically involve singular, high-stakes decisions with no equivalent alternatives. In such critical situations, the presence of arbitrariness raises serious concerns for the responsible deployment of machine learning models.
As explored earlier, the negative effects of multiplicity extend beyond the arbitrariness of just the final prediction or decision (§2.2). For instance, multiplicity in counterfactual explanations can impact the stability and validity of algorithmic recourse (Sreedharan et al., 2018; Hancox-Li, 2020; Hasan and Talbert, 2022; Leventi-Peetz and Weber, 2022; Baniecki et al., 2024; Hamman et al., 2023; Leofante et al., 2023; Kobylińska et al., 2024; Jiang et al., 2024a). The feasibility or nature of recourse might hinge on arbitrary design choices made during model development. These decisions can have real-world implications; for instance, a recourse option provided by one model may become invalid if the model is updated, invalidating previous efforts. This inconsistency raises both ethical and legal concerns (Kobylińska et al., 2024; De Toni et al., 2024).
This arbitrariness becomes even more problematic when it disproportionately impacts different individuals, particularly harming underrepresented demographics. As hinted earlier, a significant source of arbitrariness is the model’s lack of ability to learn the underlying distribution. In modern machine learning, underrepresented groups often face these information gaps, which can manifest as data scarcity due to limited historical records or a lack of understanding of cultural context for how data relates to predictions (Barocas et al., 2023; Hooker, 2021; Buolamwini and Gebru, 2018; Birhane, 2022; Geirhos et al., 2020). Research has consistently shown that such disparities exacerbate existing inequalities—whether through arbitrariness, uncertainty, or multiplicity (Farnadi et al., 2024). The resulting disproportionate harms across groups highlight the pressing need to address arbitrariness in critical domains (Gomez et al., 2024; Cooper et al., 2024; Ali et al., 2021).
Impact of Other Responsible AI Constraints on Multiplicity: We saw that multiplicity is a critical consideration in responsible AI. Interestingly, several works have also explored how multiplicity interacts with other pillars of responsible AI. For instance, studies have shown that imposing fairness constraints can inadvertently increase multiplicity (Long et al., 2024; Cavus and Biecek, 2024a). However, Long et al. (2024) argue that multiplicity stands outside the fairness-utility trade-off, meaning improvements in fairness do not have to entail increased multiplicity. They demonstrate that multiplicity can often be reduced through techniques like ensembling, while maintaining fairness. Similarly, Kulynych et al. (2023) explore the interaction between differential privacy and multiplicity, revealing that introducing privacy constraints tends to increase multiplicity. These findings underscore the complex interplay between multiplicity and other responsible AI principles, highlighting the need for further research to understand and navigate these trade-offs effectively.
6.2. Multiplicity and Homogenization
As discussed above, the extent to which arbitrariness is problematic often depends on the context, and in certain cases, it might not be a standalone concern. For instance, Creel and Hellman (2022) argue that arbitrariness in hiring decisions, by itself, is neither a legal nor moral issue. Instead, they suggest that the absence of arbitrariness across systems could lead to a different concern, creating an ‘algorithmic leviathan’, i.e., the standardization of a single outcome across an entire sector. Kleinberg and Raghavan (2021) discuss a similar concern in the form of ‘algorithmic monoculture’, which would be particularly problematic in interconnected systems, for instance, when multiple banks assess an individual’s creditworthiness, algorithmic monoculture would imply that an individual rejected from one bank would be rejected from all banks. Please refer to Table 2 for an overview of other common terms used in this literature.
This phenomenon, known as outcome homogenization, refers to the convergence of decisions due to common design choices across multiple models. In our ICA framework (§2.1), we had termed these as conventional choices. Bommasani et al. (2022) shows that outcome homogenization can occur even when different algorithms share only certain components, i.e., homogenization can occur even when only some design choices are conventional.
Term Definition
Algorithmic Blackballing (Ajunwa, 2020) ‘A worker’s lack of control over the portability of applicant data captured by automated hiring systems […] raising the specter of an algorithmically permanently excluded class of job applicants’
Algorithmic Monoculture (Kleinberg and Raghavan, 2021) ‘The notion that choices and preferences will become homogeneous in the face of algorithmic curation’
Algorithmic Leviathan (Creel and Hellman, 2022) ‘Automated decision-making systems that make uniform judgments across broad swathes of a sector.’
Outcome Homogenization (Bommasani et al., 2022) ‘The phenomenon of individuals (or groups) exclusively receiving negative outcomes from all ML models they interact with’
Generative Monoculture (Wu et al., 2024) ‘A distribution shift [towards less varied outputs] from source data (i.e., human-generated training data) to model-generated data (i.e., model outputs) for a specific task.’
Algorithmic Pluralism (Jain et al., 2024b) ‘A state of affairs in which the algorithms used for decision-making are not so pervasive and/or strict as to constitute a severe bottleneck on opportunity.’
Table 2. Various terms used in the homogenization literature.
Interestingly, introducing controlled arbitrariness or multiplicity can mitigate these risks by preventing monocultures. In contexts where arbitrariness in individual decisions is less concerning but homogenization is, controlled multiplicity becomes a desirable property (Barocas et al., 2023; Perry and Zarsky, 2015; Jain et al., 2024a). In such situations, when an intentional design choice is not possible, the developers should prefer arbitrary choices over conventional ones wherever feasible (§2.1). Despite this idea being widely recognized in the academic literature (Barocas et al., 2023; Perry and Zarsky, 2015; Jain et al., 2024a), public perception of intentional randomness in decision-making remains skeptical. A recent study by Meyer et al. (2024) indicates a strong aversion of the end users towards any form of randomization or intentional arbitrariness in automated decision-making. Therefore, fostering greater public awareness about the nuanced impacts of multiplicity is crucial before we can develop and employ potential solutions (Barocas et al., 2023).
7. Open Research Questions and Emerging Trends
Building upon our systematic survey of the literature on multiplicity, this work gives us a unique vantage point to identify and discuss several emerging trends and critical open research questions in the field,
  • Expanding the scope of multiplicity beyond predictions and explanations. One of our key motivations for broadening the definition of multiplicity (§3.1) was to incorporate multiplicity beyond predictive and explanatory contexts. Although interest in these aspects is growing, further work is needed to bring them to the forefront. We hope our work will encourage future research that fosters a deeper and more holistic understanding of multiplicity’s broader impact.
  • Cost-effective enumeration of Rashomon sets. A major challenge in auditing multiplicity lies in the resource-intensive nature of training multiple models to enumerate the Rashomon set. While we discussed several works that improve the efficiency of this enumeration (§3.2), the need for further research in this direction remains pressing.
  • Mathematical foundations of multiplicity. Establishing stronger mathematical foundations of multiplicity, for instance, our focus on formalizing the distinction between multiplicity, uncertainty, and bias-variance decomposition (§4), is essential. Fundamental work on the Rashomon effect in machine learning is less represented (only $12.5\%$; see Figure 1), highlighting the opportunities for future work on frameworks that rigorously define and explore multiplicity.
  • Multiplicity and its interaction with responsible AI. Given the conversation of arbitrariness as a concern of responsible model development (§6.1), its interaction with other pillars of responsible AI is warranted. Future research on frameworks that address multiplicity within the broader landscape of responsible AI deployment is needed.
  • Interdisciplinary perspective on multiplicity. While substantial work has been done on measuring and mitigating multiplicity, many studies do not engage further with the concerns of multiplicity in real-world settings. As discussed, arbitrariness can be both a force of good (§6.2) or a cause of harm (§6.1), depending on the context. Our systematic review shows that only $47.5\%$ of works explicitly engage with responsible AI concerns (see Figure 1), which we believe to be low given the field’s relevance to the responsible deployment of machine learning models. Future collaborative efforts across disciplines are crucial to truly grasp the real-world consequences of multiplicity.
  • Multiplicity in the era of LLMs. As machine learning models continue to scale, particularly in the era of large language models (LLMs), new challenges emerge. For instance, we see increasing attention given to the concerns of monoculture and homogenization (§6.2). Even the evaluation of multiplicity becomes increasingly complex, as training multiple models is often infeasible at this scale. Additionally, we see another dimension of multiplicity in LLMs–decisions multiplicity due to changing prompts–that requires deeper examination (Lu et al., 2022; Alzahrani et al., 2024; Sclar et al., 2023; Voronov et al., 2024).
8. Conclusion
In this work, we systemized the existing knowledge on multiplicity, uncovering interesting trends and insights. One limitation of our study is the evolving terminology within the field–terms such as “multiplicity” and “Rashomon sets” have gained prominence only recently. As a result, our survey may have missed relevant works that did not explicitly use this terminology. Despite this, our efforts to formalize key discussions–the language around developer choices, definitions, and the distinction between multiplicity, uncertainty, and variance–represent a crucial step toward unifying the field. We also explored broad trends related to the real-world impacts of multiplicity, building from existing literature and highlighting overarching themes that extend beyond their originally studied contexts. We hope our work provides a platform that is both accessible to newcomers and valuable to experts, fostering further research in multiplicity.
Acknowledgements.
Funding support for project activities has been partially provided by Canada CIFAR AI Chair, Google award, FRQNT and NSERC Discovery Grants program. We also express our gratitude to Compute Canada for their support in providing facilities for our evaluations.
Appendix A Systematic Literature Review
We conducted a systematic literature review to collect and analyze works related to multiplicity. Our process involved the following steps:
  1. (1)
    Paper Collection: We indexed various online repositories using a set of predefined search terms. Details of the search are provided in §A.1. After filtering for duplicates across various search terms and repositories, we retained a total of 339 papers.
  2. (2)
    Manual Filtering: We manually reviewed the remaining papers to exclude those not relevant to our study. The exact inclusion criteria are noted in §A.2. This manual filtering reduced the dataset to 80 papers.
  3. (3)
    Manual Tagging: Finally, each paper was assigned tags to identify and emphasize the specific aspects of multiplicity it addressed. The set of tags used and categorization rules are discussed in §A.3.
After filtering and tagging the papers, we analyzed various statistics, as shown in the main text. The following sections provide a detailed account of each step of the systematic review process.
A.1. Paper Collection
We indexed and collected papers using the search terms ‘rashomon’, ‘model multiplicity’, and ‘set of good models’ from two primary sources: DBLP [^1] https://dblp.org/ and the ACM Digital Library [^2] https://dl.acm.org/. To ensure comprehensive coverage, we used the extended search in ‘The ACM Guide to Computing Literature’ within the ACM Digital Library. The end date for our search was 31 Dec 2024, and no restrictions were put on the starting date. Together, DBLP and the ACM Digital Library provide coverage of major machine learning conferences (e.g., NeurIPS, ICML, ICLR, AAAI, ACL, NAACL), leading venues focused on responsible AI (e.g., FAccT, AIES, EAAMO), and archival repositories such as arXiv. The exact number of papers collected from each source and search query is detailed in Table 3.
Search Term Number of Papers
DBLP ACM Digital Library Combined (Duplicates Removed)
‘rashomon’ 34 156 178
‘model multiplicity’ 66 95 153
‘set of good models’ 11 28 38
Total Number of Papers (Duplicates Removed):    339
Total Number of Papers (After Manual Filtering):      80
Table 3. Paper collection statistics.
A.2. Manual Filtering
After removing duplicates, we had 339 papers in our dataset. Each paper was then manually reviewed to determine its relevance to our survey. Our focus was on works that deeply engaged with the Rashomon effect and multiplicity within the context of machine learning. Thus, papers examining the Rashomon effect outside of machine learning, investigating multiplicity in other domains, or briefly mentioning the Rashomon effect without further exploring it, were excluded. The detailed inclusion and exclusion criteria are outlined in Table 4.
Inclusion Criteria Exclusion Criteria
- Papers whose central contribution was deeply intertwined with the Rashomon effect or multiplicity. - Papers that mentioned the Rashomon effect, but did not engage with it in their main contributions.
- Position papers and surveys that devoted a considerable amount of space to the Rashomon effect or multiplicity (We used one complete section worth of space as our threshold). - Papers exploring the Rashomon effect or model multiplicity in domains outside machine learning. The term ‘model’ in these cases did not refer to machine learning models.
- Commentaries, extended abstracts, tutorials, complete conference proceedings, and other indexed documents that are not considered research papers.
- Duplicates for papers which exist in various repositories under different names, and hence could not be filtered automatically.
Table 4. Inclusion and exclusion criteria for manual filtering.
A.3. Manual Tagging
Finally, after manual filtering, we were left with 80 papers. Each paper was then tagged based on their contributions in the context of multiplicity in machine learning. The tags were created by the authors after a preliminary review of the papers during the manual filtering step. Strict rules were defined for each tag, as shown in Table 5. Every paper was assigned all applicable tags through this manual tagging process. These 80 tagged papers formed the basis of the statistics presented in Fig 1 of the main text.
Tag Rule
Rashomon Effect in ML Papers that provide fundamental insights into the Rashomon effect in ML, including its causes, forms of manifestation, broader characteristics, etc.
Rashomon Set Exploration Papers that focus on the problem of enumerating the Rashomon set, including papers which highlight properties of the Rashomon set. This does not include papers that already assume access to the Rashomon set.
Better Models and Ensembles Papers that use the Rashomon set to find better models, or combine multiple models from the Rashomon set to improve certain objectives. Thus, any work that takes advantage of the flexibility the Rashomon set provides.
Hacking Metrics Papers that highlight the negatives of underspecification, using the Rashomon effect to hack existing metrics, checks, or regulations.
Responsible AI Papers that explicitly engage with the harms or benefits of the Rashomon effect in the context of Responsible AI. Papers whose central contribution is related to Responsible AI, but do not engage with that aspect of their impact, were not included here. For example, papers tackling explanation multiplicity strictly from a technical point of view were not given this tag.
Application Papers whose central contribution is limited to a particular application.
Survey Position papers and surveys that either provide an overarching discussion on multiplicity or place multiplicity in a broader context.
Predictive Multiplicity Papers that focus on predictive multiplicity, i.e., when models in the Rashomon set have varying predictions.
Explanation Multiplicity Papers that focus on explanation multiplicity, i.e., when models in the Rashomon set have varying explanations.
Other Multiplicity Papers that focus on any other form of multiplicity beyond predictive or explanation multiplicity. This can include fairness multiplicity, OOD robustness multiplicity, model complexity multiplicity, feature interaction multiplicity, etc.
Table 5. Rules for each tag in the manual tagging step.